# A Unified Framework for Model Editing

### v1

@@ Abstract
Model editing is a growing area focused on updating the knowledge embedded within models. Among the various methodologies, ROME and MEMIT stand out as leading "locate-and-edit" model editing techniques. While MEMIT enables batched editing of memories, ROME is limited to changing one fact at a time. This paper introduces a unifying framework that brings ROME and MEMIT under a single conceptual umbrella, optimizing for the same goal, which we call the preservation-memorization objective. This objective aims to preserve the representations of certain selected vectors while memorizing the representations of new factual information. Specifically, ROME optimizes this objective using an equality constraint, whereas MEMIT employs a more flexible least-square constraint. In addition to making batched edits, MEMIT also edits the model at multiple layers. We disentangle the distribution of edits to multiple layers from the optimization objective of MEMIT and show that these edit-distribution algorithms should be considered separate entities worthy of their own line of research.
Finally, we present EMMET - an Equality-constrained Mass Model Editing algorithm for Transformers, a new batched memory-editing algorithm. With EMMET, we present a closed form solution for the equality-constrained version of the preservation-memorization objective. We show that EMMET is able to perform batched-edits on par with MEMIT up to a batch-size of 256 and discuss the challenges in stabilizing EMMET. By articulating the "locate-and-edit" model editing algorithms under a simple conceptual framework of preservation-memorization, we aim to bridge the gap between intuition and mathematics and hope to simplify the journey for future researchers in model editing.
@@ 1 Introduction
As new facts emerge constantly, it’s crucial to keep models up-to-date with the latest knowledge. Model editing gives us the ability to edit facts stored inside a model as well as update incorrectly stored facts. Model editing methods can be broadly classified into two types - methods that add information in-context (Mitchell et al., 2022; Zhong et al., 2023; Cohen et al., 2023), and methods that modify the parameters of underlying model (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a, b; Tan et al., 2023). Various model editing techniques have been proposed in the past that tackle this problem in different ways. Dai et al. (2021) first identify knowledge containing neurons in a model using integrated gradients (Sundararajan et al., 2017) and then modify the selected neurons to edit facts in a model. This method is not scalable with increasing model sizes as it requires us to find activations for each neuron in the model. De Cao et al. (2021) and Mitchell et al. (2021) train a hypernetwork (Chauhan et al., 2023) that generates the new weights of the model being edited. While these methods have been optimized to scale with a square-root dependence on the size of the edited model, it still requires training of additional editing models dependent on each source model being edited. Other methods add the most relevant updated knowledge in context (Mitchell et al., 2022; Cohen et al., 2023; Zhong et al., 2023). While such methods provide a viable alternative to model editing, in this paper, we focus on parameter-modifying model editing methods.
Figure 1: Figure shows a diagrammatic representation of a transformer layer. The layer being edited by ROME, MEMIT and EMMET is the project weight matrix inside the MLP layer ($W_{proj}$), as highlighted in the figure.
We search for model editing methods that can insert new knowledge into the model weights, while being universally applicable to any model. This goal requires a better understanding of the knowledge storing mechanisms in LLMs and in the process of solving this problem, we also hope to make these models more interpretable. So far, the most promising model editing methods that universally modify model weights are ROME (Rank-One Model Editing) (Meng et al., 2022a; Gupta and Anumanchipalli, 2024) and MEMIT (Mass Editing Memory in Transformer) (Meng et al., 2022b). These methods directly update specific "knowledge-containing" parts of the model without requiring the need to train additional models and can be applied to any transformer based large language model (LLMs). MEMIT also presents the only existing non-hypernetwork based parameter modifying model editing method that allows for batched edits.
@@
In this paper, we present a unifying conceptual framework for ROME and MEMIT and show that both methods optimize the same objective. We call this the preservation-memorization objective of model editing, where new knowledge is injected or memorized such that representations of certain vectors are preserved through the editing process. We show that ROME optimizes an equality-constrained version of the objective whereas MEMIT optimizes a more relaxed least-squares version of the objective, which allows for a simple closed-form solution for making batched edits. We then highlight that MEMIT consists of two separate steps - an optimization objective and an algorithm that distributes the edits into multiple layers. We show that while MEMIT optimizes a more relaxed objective, the power of MEMIT in many cases comes from these edit-distribution algorithms. We disentangle the edit-distribution algorithms from the optimization objective of MEMIT and show that these edit-distribution algorithms are universally applicable. Our experiments show that while MEMIT’s edit-distribution algorithm helps editing with large batches for GPT2-XL (Radford et al., 2019) and GPT-J (Wang and Komatsuzaki, 2021), it hurts model editing performance when used with Llama-2-7b (Touvron et al., 2023). With this, we advocate for further research into the edit-distribution algorithms and viewing them as separate entities from the optimization objectives.
Finally, we present the closed form solution for making batched edits with the equality-constrained preservation-memorization objective in the form of EMMET - an Equality-constrained Mass Model Editing algorithm for Transformers. EMMET is a new batched-editing algorithm that provides symmetry between the two objectives of "locate-and-edit" class of algorithms and show that batched edits can be made using both objectives. While EMMET does batched editing under an equality constraint, enforcing a "theoretically" more accurate editing of memories, in practice, we find that EMMET performs on-par with MEMIT for batch sizes only up to 256. We believe the reason for this is the hard equality constraints enforced in the optimization objective of EMMET. While EMMET performance does not match MEMIT for larger batch sizes, it can still be used for sequential model editing with smaller batch sizes (Yao et al., 2023; Gupta et al., 2024). The code for EMMET can be found here.
@@ 2 Background
Model editing, as traditionally defined [^1] We call this the traditional definition of model editing since non-parameter modifying methods that add new knowledge in context are now also considered part of the model editing domain., is the task of adjusting the model’s parameters such that a new fact is incorporated into the model weights without influencing its behavior on other samples (Yao et al., 2023). Facts are usually represented in a key-value format when presented for editing. As an example, let us say we are adding a new fact into the model - "The president of USA is John Cena". Then this fact is represented by a pair of key-value vectors ($k_{e}$, $v_{e}$), where $k_{e}$ is the vector representation of the phrase - "The president of USA is" that maximally corresponds to retrieval of the fact, and $v_{e}$ is the vector representation of the output of the key vector ($k_{e}$) at the layer being edited such that "John Cena" is produced as output from the model. This is pictorially represented in step-2 in Figure 2. For a more detailed explanation of creation of key-value vectors, we refer the readers to (Meng et al., 2022a).
While early model editing methods made singular but accurate edits, more recently, focus has shifted to scaling up model editing methods. Scaling of model editing can happen in two ways - either through sequential editing (Yao et al., 2023; Gupta et al., 2024) or batched editing (Meng et al., 2022b). In sequential editing, edits are made to the model sequentially whereas in batched editing, multiple facts are inserted into the model with a single gradient update. MEMIT (Meng et al., 2022b) is one such batched editing algorithm. In this paper, we present EMMET, which is a new batched editing algorithm that inserts multiple facts into the model with one gradient update.
The success of model editing is measured using standard model editing metrics following the work of (Meng et al., 2022a). We use the following five model editing metrics:
@@
  • Paraphrase Score (PS) represents the generalization ability of model under an edit. It is measured as percentage of edits where $P(\text{new fact})>P(\text{old fact})$ under paraphrases of the query prompt.
  • Neighborhood Score (NS) represents locality of model editing. In other words, it measures if editing of a fact effects other facts stored inside a model. NS represents the percentage of facts in the neighborhood of the edited fact that remain unaltered post-edit.
  • Generation Entropy (GE) represents the fluency of a model post edit. It is calculated by measuring the weighted average of bi-gram and tri-gram entropies of text generated by an edited model. This quantity drops if the generated text is repetitive, a common failure case of model editing.
  • Score (S) is a quantify defined by (Meng et al., 2022a) to represent a combination of edit success, generalization and locality. It is the harmonic mean of ES, PS and NS.
@@ 3 Preservation-Memorization : A Unifying Framework for ROME and MEMIT
ROME and MEMIT are two of the most popular model editing methods. While ROME allows only singular edits, MEMIT allows for editing multiple memories in a model at the same time, thus providing a scalable solution to model editing. Both algorithms base their work on viewing the weights of the feed-forward layer in a transformer as linear associative memories (Kohonen, 1972; Anderson, 1972). Under this paradigm, linear operations in a transformer (feed-forward layers) are viewed as a key-value store for information. In this section, we re-introduce both ROME and MEMIT in a new light - a unifying conceptual framework of the preservation-memorization objective.
@@
Let $W$ represent the weights of a feed-forward layer in a transformer deemed worthy of editing [^2] These layers are found by causal tracing methods (Meng et al., 2022a, b) to modify existing knowledge, and let $k$ be a key-vector representative of a fact that we are either editing or preserving, and is the input vector to $W$. The layers being edited are shown in an expanded diagram of a transformer layer in Figure 1. In the model editing process, the weights of an intermediate layer of the model are changed from $W_{0}$ to $\hat{W}$, where $k_{0}$ is used to indicate a key-vector representing facts we want to preserve from the original model, and $k_{e}$ being key-vectors representing facts we want to insert into the model. Let $v_{e}$ be the desired output at the layer being edited corresponding to input $k_{e}$ such that the correct fact is recalled by the model when finally generating text. For more details on how we find $k_{e}$ and $v_{e}$, we refer the reader to Meng et al. (2022a).
@@
Our objective is then to preserve the representations of selected input vectors before and after editing, or in other words, minimize the error between $W_{0}k_{0}$ and $\hat{W}k_{0}$, while forcing the output representation of the vector $k_{e}$ to be $v_{e}$, or in other words - memorizing the fact represented by ($k_{e}$, $v_{e}$). This process is shown pictorially in Figure 2. This problem can be posed in two different ways. In ROME-style, this objective of model editing is optimized by the following equation:
@@
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 4.0pt\underbrace{\left\|\hat{W}K_{0}-W_{0}K_{0}\right\|}_{\text{preservation}}\hskip 4.0pt\text{ such that }\underbrace{\hat{W}k_{e}=v_{e}}_{\text{memorization}}$ (1)
where $K_{0}=[k^{0}_{1}\hskip 4.0pt|k^{0}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptk^{0}_{N}]$ is a matrix containing all the vectors whose representations we want to preserve in a row. The first term in the above equation represents preservation of representations of selected vectors in $K_{0}$ while the second term enforces memorization of the fact represented by ($k_{e}$, $v_{e}$) using an equality constraint.
[…]
The solution for the above objective can be written as:
$\hat{W}=W_{0}+\Delta\hskip 10.0pt\text{where}\hskip 10.0pt\Delta=(v_{e}-W_{0}k_{e})\frac{k_{e}^{T}C_{0}^{-1}}{k_{e}^{T}C_{0}^{-1}k_{e}}$ (2)
@@
where $K_{E}=[k^{e}_{1}\hskip 4.0pt|k^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptk^{e}_{E}]$ is a matrix containing a row of vectors representing the edits we are making and $V_{E}=[v^{e}_{1}\hskip 4.0pt|v^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptv^{e}_{E}]$ represents their target representations. We again see that the first term in the above equation preserves the representation of selected vectors while the second terms forces memorization of facts represented by $(K_{E},V_{E})$ using a least square constraint.
@@
The above optimization objective aims to modify the output representations of vectors in $K_{E}$ to $V_{E}$ by minimizing the least square error between them instead of requiring them to be equal with an equality constraint. This is the major difference between the objectives of ROME and MEMIT, where ROME poses the memorization part of the objective as an equality constraint whereas MEMIT relaxes the equality constraint to a least-square objective. This allows Meng et al. (2022b) to find a closed form solution for making $E$ edits to the model in a single update, represented by the matrix $K_{E}$. The solution for the MEMIT objective is:
$\hat{W}=W_{0}+\Delta\hskip 10.0pt\text{where}\hskip 10.0pt\Delta=\big{(}V_{E}-W_{0}K_{E}\big{)}K_{E}^{T}\big{(}\lambda C_{0}+K_{E}K_{E}^{T}\big{)}^{-1}$ (4)
We deliberately write the first term in both solutions in a similar form. The first term in $\Delta$ represents the residual error (represented by $R$) of the new associations ($K_{E},V_{E}$) when evaluated on the old weights. $R\triangleq v_{e}-W_{0}k_{e}$ is a vector in case of ROME since we are only able to make singular edits, whereas $R\triangleq V_{E}-W_{0}K_{E}$ is a matrix for MEMIT consisting of a row of vectors corresponding to each edit in the batch.
To summarize, ROME and MEMIT can be seen as two realizations of the preservation-memorization (PM) objective of model editing, where ROME enforces memorization using an equality constraint whereas MEMIT enforces memorization as a least square objective. The least-square constraint in MEMIT allows to reach a closed form solution for batch updates.
@@
(d) Generation Entropy (GE)
(e) Score (S)
Figure 3: Performance comparison of model editing using MEMIT when editing just one layer against multiple layers using the MEMIT edit-distribution algorithm on the CounterFact dataset. The multi-layer edits are represented by dotted lines whereas single layer edits are represented by solid lines.
Edit-Distribution Algorithms.

### v2

@@ Abstract
We introduce a unifying framework that brings two leading "locate-and-edit" model editing techniques – ROME and MEMIT – under a single conceptual umbrella, optimizing for the same goal, which we call the preservation-memorization objective. ROME uses an equality constraint to perform one edit at a time, whereas MEMIT employs a more flexible least-square constraint that allows for batched edits. Following the preservation-memorization objective, we present Equality-constrained Mass Model Editing algorithm for Transformers or EMMET, a new batched memory-editing algorithm that uses a closed-form solution for the equality-constrained version of the preservation-memorization objective. EMMET is a batched-version of ROME and is able to perform batched-edits up to a batch-size of 10,000 with very similar performance to MEMIT across multiple dimensions. With EMMET, we unify and achieve symmetry within the "locate-and-edit" algorithms, allowing batched-editing using both objectives.
A Unified Framework for Model Editing
Akshat Gupta, Dev Sajnani, Gopala Anumanchipalli UC Berkeley {akshat.gupta, sajnanidev, gopala}@berkeley.edu
@@ 1 Introduction
As new facts emerge constantly, it’s crucial to keep models up-to-date with the latest knowledge. Model editing Yao et al. (2023) gives us the ability to edit facts stored inside a model as well as update incorrectly stored facts. In this paper, we focus on two popular parameter modifying model editing methods which infuse knowledge within models without needing an additional hypernetwork Chauhan et al. (2023). These methods are ROME (Rank-One Model Editing) (Meng et al., 2022a) and MEMIT (Mass Editing Memory in Transformer) (Meng et al., 2022b). These methods directly update specific "knowledge-containing" parts of the model without requiring the need to train additional models and can be applied to any transformer based large language model (LLMs). MEMIT also uniquely allows for batched edits (appendix A.1).
@@
In this paper, we present a unifying conceptual framework for ROME and MEMIT and show that both methods optimize the same objective. We call this the preservation-memorization objective of model editing, where new knowledge is injected or memorized such that representations of certain vectors are preserved through the editing process. We show that ROME optimizes an equality-constrained version of the objective whereas MEMIT optimizes a more relaxed least-squares version of the objective, which allows for a simple closed-form solution for making batched edits. We then highlight that MEMIT consists of two separate steps - an optimization objective and an algorithm that distributes the edits into multiple layers. The power of MEMIT in many cases comes from these edit-distribution algorithms.
Finally, we present a closed-form solution for making batched edits with the equality-constraint under the preservation-memorization objective in the form of EMMET - an Equality-constrained Mass Model Editing algorithm for Transformers. With EMMET, batched edits can be performed for batch sizes upto 10,000 with its performance matching MEMIT across multiple dimensions. We evaluate EMMET on three models - GPT2-XL Radford et al. (2019), GPT-J Wang and Komatsuzaki (2021) and Llama-2-7b Touvron et al. (2023) on standard model editing datasets - CounterFact Meng et al. (2022a, b) and zsRE Levy et al. (2017). The code for EMMET can be found here [^1] https://github.com/myanonymousrepo/unified_model_editing.
The main contributions of our paper are:
  • We unify two popular model editing techniques (ROME and MEMIT) under the same conceptual framework called the preservation-memorization.
  • We disentangle the MEMIT objective from the MEMIT algorithm which distributes edits within multiple layers. We hope this sparks further research in edit-distribution algorithms.
  • We present a closed-form solution to equality-constrained memorization in the form of EMMET, a batched version of ROME. EMMET is a new batched-editing algorithm that achieves symmetry between the two objectives of "locate-and-edit" class of algorithms and shows that batched edits can be made using both objectives.
@@ 2 Background
Facts for model editing are usually represented in a key-value format where the key vector has maximal correspondence to retrieval of a fact and the value vector enables us to get the target output after editing (Meng et al., 2022a; Geva et al., 2020). As an example, let us say we are editing a new fact into the model - "The president of USA is John Cena". In this fact, $k_{e}$ is the vector representation of the phrase - "The president of USA is," and $v_{e}$ is the vector representation of the output at the layer being edited such that "John Cena" is produced as output at the final layer of the model. This is pictorially represented in step 2 in Figure 1. For a more detailed explanation of the creation of key-value vectors, we refer readers to (Meng et al., 2022a).
The success of model editing is measured using standard model editing metrics (Meng et al., 2022a; Yao et al., 2023) described below:
@@
  • Paraphrase Score (PS) represents the generalization ability of model under an edit. It is measured as the percentage of edits where $P(\text{new fact})>P(\text{old fact})$ under paraphrases of the query prompt.
  • Neighborhood Score (NS) represents locality of model editing. In other words, it measures if editing of a fact affects other facts stored inside a model. NS represents the percentage of facts in the neighborhood of the edited fact that remain unaltered post-edit.
  • Generation Entropy (GE) represents the fluency of a model post edit. It is calculated by measuring the weighted average of bi-gram and tri-gram entropies of text generated by an edited model. This quantity drops if the generated text is repetitive, a common failure case of model editing (Meng et al., 2022a; Gupta and Anumanchipalli, 2024).
  • Score (S) is a quantify defined by (Meng et al., 2022a) to represent a combination of edit success, generalization and locality. It is the harmonic mean of ES, PS, and NS.
Figure 2: Figure shows a diagrammatic representation of a transformer layer. The layer being edited by ROME, MEMIT and EMMET is the projection weight matrix inside the MLP layer ($W_{proj}$).
@@ 3 Preservation-Memorization : A Unifying Framework for ROME and MEMIT
Both ROME and MEMIT base their work on viewing the weights of the feed-forward layer in a transformer as linear associative memories (Kohonen, 1972; Anderson, 1972). Under this paradigm, linear operations in a transformer (feed-forward layers) are viewed as a key-value store for information. In this section, we re-introduce both ROME and MEMIT in a new light - a unifying conceptual framework of the preservation-memorization objective.
@@
Let $W$ represent the weights of the feed-forward layer we want to edit [^2] These layers are found by causal tracing methods (Meng et al., 2022a, b), and let $k$ be a key-vector representative of a fact that we are either editing or preserving, and is the input vector to $W$. The layers being edited are shown in an expanded diagram of a transformer layer Vaswani et al. (2017) in Figure 2. The output of a single transformer layer for a general decoder-only LLM can be written as:
$a^{l}$ $={LN}(\texttt{Att}(h^{l-1})+h^{l-1})$ (1)
$m^{l}$ $=W^{l}_{proj}\left({NL}(W^{l}_{fc}a^{l}+b^{l}_{fc})\right)+b^{l}_{proj}$ (2)
$h^{l}$ $={LN}(m^{l}+a^{l})$ (3)
Here, Att is the multi-head attention function, NL refers to the non-linearity used in the MLP module of the model and LN refers to layer normalization. The keys are outputs of the first linear layer, or $k^{l}=\texttt{NL}(a^{l}W^{l}_{fc}+b^{l}_{fc})$, whereas $v^{l}=m^{l}$. A detailed explanation on creation of key-vectors and value-vectors is given in Appendix A.3.
In the model editing process, the weights of an intermediate layer of the model are changed from $W_{0}$ to $\hat{W}$, where $k_{0}$ is used to indicate a key-vector representing facts we want to preserve from the original model, and $k_{e}$ being key-vectors representing facts we want to insert into the model. Let $v_{e}$ be the desired output at the layer being edited corresponding to input $k_{e}$ such that the correct fact is recalled by the model when finally generating text.
@@
Our objective is then to preserve the representations of selected input vectors before and after editing, or in other words, minimize the error between $W_{0}k_{0}$ and $\hat{W}k_{0}$, while forcing the output representation of the vector $k_{e}$ to be $v_{e}$, or in other words - memorizing the fact represented by ($k_{e}$, $v_{e}$). This process is shown pictorially in Figure 1.
In ROME-style, this objective of model editing is optimized by the following equation:
@@
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 4.0pt\underbrace{\left\|\hat{W}K_{0}-W_{0}K_{0}\right\|}_{\text{preservation}}\hskip 4.0pt\text{ s.t. }\underbrace{\hat{W}k_{e}=v_{e}}_{\text{memorization}}$ (4)
where $K_{0}=[k^{0}_{1}\hskip 4.0pt|k^{0}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptk^{0}_{N}]$ is a matrix containing all the vectors whose representations we want to preserve in a row.
We call this the preservation-memorization objective of model editing because it allows us to retain existing knowledge or skills of a model by keeping the same representations of selected key-vectors before and after editing, while memorizing a new fact $k_{e}$, whose representation are forced to be $v_{e}$, where $v_{e}$ is by definition the output representation for $k_{e}$ that generates the target answer at final layer.
The solution for ROME can then be written as:
$\hat{W}$ $=W_{0}+\Delta\hskip 10.0pt\text{where}\hskip 10.0pt$ (5)
$\Delta$ $=(v_{e}-W_{0}k_{e})\frac{k_{e}^{T}C_{0}^{-1}}{k_{e}^{T}C_{0}^{-1}k_{e}}$ (6)
@@
where $K_{E}=[k^{e}_{1}\hskip 4.0pt|k^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptk^{e}_{E}]$ is a matrix containing a row of vectors representing the edits we are making and $V_{E}=[v^{e}_{1}\hskip 4.0pt|v^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptv^{e}_{E}]$ represents their target representations.
@@
The above optimization objective aims to modify the output representations of vectors in $K_{E}$ to $V_{E}$ by minimizing the least square error between them instead of requiring them to be equal with an equality constraint. This is the major difference between the objectives of ROME and MEMIT, where ROME poses the memorization part of the objective as an equality constraint whereas MEMIT relaxes the equality constraint to a least-square objective. This allows Meng et al. (2022b) to find a closed-form solution for making $E$ edits to the model in a single update, represented by the matrix $K_{E}$. The solution for the MEMIT objective is:
$\hat{W}$ $=W_{0}+\Delta\hskip 10.0pt\text{where}\hskip 10.0pt$ (8)
$\Delta$ $=\big{(}V_{E}-W_{0}K_{E}\big{)}K_{E}^{T}\big{(}\lambda C_{0}+K_{E}K_{E}^{T}\big{)}^{-1}$
We deliberately write the first term in both solutions in a similar form. The first term in $\Delta$ represents the residual error (represented by $R$) of the new associations ($K_{E},V_{E}$) when evaluated on the old weights $W_{0}$. $R\triangleq v_{e}-W_{0}k_{e}$ is a vector in case of ROME since we are only able to make singular edits, whereas $R\triangleq V_{E}-W_{0}K_{E}$ is a matrix for MEMIT consisting of a row of vectors corresponding to each edit in the batch.
To summarize, in this section we show that ROME and MEMIT can be seen as two realizations of the preservation-memorization (PM) objective of model editing, where ROME enforces memorization using an equality constraint whereas MEMIT enforces memorization as a least square objective. The least-square constraint in MEMIT allows to reach a closed form solution for batch updates.
@@
Figure 3: Performance comparison of model editing using MEMIT when editing just one layer against multiple layers using the MEMIT edit-distribution algorithm on the CounterFact dataset.
4 Edit-Distribution Algorithms